531.05 J of Work are done on a system in a process that decreases the systems Thermal Energy by 569.73 J.

How much energy is transferred to(+ answer) or from the system (- answer) as heat?

Your answer should have 2 digits after the decimal.  Rounding rules apply.
Answer #1

We know that:

∆U = W + Q

Hense, Q = 569.73 - 531.05 = 38.68 J

Hense, 38.68 J of heat is transferred to the system.